OpenStudy (anonymous):

solve for Y: sqrt(y) + sqrt(y-5) = 5

OpenStudy (tkhunny):

FIRST, think about the Domain. \(\sqrt{y}\) DEMANDS that \(y \ge 0\) \(\sqrt{y-5}\) DEMANDS that \(y \ge 5\) Having said that, WRITE DOWN \(y \ge 5\), then isolate one radical and square both sides.

OpenStudy (anonymous):

Thank you!